Solve each equation.[please help with this 3 that i dont get]

8 – 3x = –1
log381 = x


log x – log 3 = 2

Answer by Alan3354(21583)

Solve each equation.[please help with this 3 that i dont get]
8 – 3x = –1
-3x = -9
x = 3
----------
log381 = x
log(381) = 2.5809... (via calculator)
--------------------
log x – log 3 = 2
log(x/3) = 2
x/3 = 100
x = 300